Thanks for the help!
Pete is three years older than his brother Al. In eight years, nine times pete's age will be equal to ten times Al's age. What is Al's present age?

Found 2 solutions by Nate, checkley71:
Answer by Nate(3500)

Equations:
P = A + 3
9(P + 8) = 10A
..................
Combine:
-9P = -9A - 27
+
9P + 72 = 10A
=
72 = A - 27
Al is 99 years old, and Pete is 102 years old.

Answer by checkley71(8403)

P+3=A & 9(P+11)=10A OR P=A-3 THUS
9(A+8)=10A OR 9A+72=10A OR A=72 THEN P=72-3 OR 69 PROOF
9(69+11)=10*72 OR 9*80=720 OR 720=720